A capacitor is in series with a resistor and is charged to a potential of 100 V. When discharging the capacitor it takes 1.2 s for the potential across the capacitor to drop to 67 V. How long does it take the potential difference across the capacitor to drop to 30 V? (to 2 s.f and in s)
1
Expert's answer
2020-05-18T10:11:57-0400

We have given that a capacitor is in series with a resistor and is charged to a potential of "100V".

Thus,"V_0 =100V" ,where "V_0" is the potential difference or emf across the capacitor.

Let "C" is the capacitance of the given capacitor and "R" is the resistance of the given resistor.

Therefore, time constant is given by "\tau =RC" .

Now, for discharging we know that voltage across the capacitor drops exponentially, which is given by

"V=V_0 e^{-\frac{t}{RC}}\\ \implies V=V_0 e^{-\frac{t}{\tau}} \hspace{1cm}(\clubs)"

Now, we have also given that potential drops across the capacitor in "t=1.2s" is "V=67 \:Volt"

Thus from "(\clubs)" we get,


"67=100e^{-\frac{1.2}{\tau}}\\
\implies ln \bigg(\frac{67}{100}\bigg)=-\frac{1.2}{\tau}\\
\implies \tau=-\frac{1.2}{ ln \bigg(\frac{67}{100}\bigg)} \hspace{1cm}(\clubs \clubs)"

Thus, on putting "(\clubs \clubs)" in "(\clubs)" we get,


"V=V_0 e^{\frac{t}{\frac{1.2}{ ln \big(\frac{67}{100}\big)}}}\\
\implies V=V_0 e^{\frac{t }{1.2} ln \big(\frac{67}{100}\big)}\\
\implies V=V_0e^{ ln \big(\frac{67}{100}\big)^{\frac{t}{1.2}}}\\
\implies V=V_0 \big(\frac{67}{100}\big)^{\frac{t}{1.2}} \hspace{1cm} (\because e^{lnx}=x) \:(\spades)"

Now from equation "(\spades)" we can calculate at "t=t_1" at where potential across the capacitor is "30V" ,


"30=100 \big(\frac{67}{100}\big)^{\frac{t_1}{1.2}}\\
\implies \frac{3}{10}=\big(\frac{67}{100}\big)^{\frac{t_1}{1.2}}\\
\implies ln \big(\frac{3}{10}\big)=\frac{t_1}{1.2}ln \big(\frac{67}{100}\big)\\
\implies t_1=1.2 \frac{ln(0.3)}{ln(0.67)}\\
\implies t_1=3.6s"

Hence, required time will be "3.6s" to reach "30V" potential across the given capacitor.